What is the concetration (mg/L) of 4.5 times 10 to the -4g of resaurin dissolved in distilled water to a volume of 15.0ml, then serially diluted four times using twofold serial dilutions?
1
Expert's answer
2019-05-24T01:59:13-0400

Initial concentration of resazurin is

"c=\frac{m}{V}=\frac{4.5\cdot 10^{-4}g}{15.0 mL}=\frac{0.45 mg}{15.0 mL}= 0.03 \frac{mg}{mL}"

After each twofold dilution concentration decreases two times. Thus after four twofold dilutions concentration decreases 24=16 times.

Final concentration of resazurin is "\frac{0.03\frac{mg}{mL}}{16}=0.001875\frac{mg}{mL}=1.875 \times 10^{-3} \frac{mg}{mL}"